OpenStudy (anonymous):

the vertex is at(-5,1), and the focus is at(2,1). write the equation and graph.

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

distance between vertex & focus= \[\sqrt{(-5-2)^{2}+(1-1)^{2}}= 7\] equation of the axis of the parabola- \[y=1\] so, equation of directrix would be- x = k (k is arbitrary constant to be determined) now, since the directrix is equidistant from vertex as focus, \[(-5-k)/\sqrt{1^{2}}=7\] so, k=-12 so the directrix is x+12=0 so the equaton of the parabola is \[(x+12)/\sqrt{1^{2}}=\sqrt{(x-2)^{2}+(y-1)^{2}}\]

OpenStudy (dumbcow):

equation: \[x = \frac{1}{28}(y-1)^{2}-5\]